\vspace{6mm}

\begin{thebibliography}{75}

\bibitem{Ahmed Fiorenza Formica at all}
{\bf I. Ahmed, A. Fiorenza, M.R. Formica, A. Gogatishvili, J.M. Rakotoson.}
{\it Some new results related to Lorentz G-Gamma spaces and interpolation.}
J. Math. Anal. Appl., 483, {\bf 2}, (2020), to appear.

\bibitem{anatriellofiojmaa2015}
{\bf G.~Anatriello} and {\bf A.~Fiorenza.} {\it Fully measurable
grand Lebesgue spaces}. J. Math. Anal. Appl. \textbf{422} (2015),
no.~2, 783--797.

\bibitem{anatrielloformicaricmat2016}
{\bf G.~Anatriello} and {\bf M.~R.~Formica.} {\it Weighted fully
measurable grand Lebesgue spaces and the maximal theorem}. Ric. Mat.
\textbf{65} (2016), no.~1, 221--233.

\bibitem{Bloom}
{\bf Steven Bloom and Gary Sampson.} {\it Weighted Spherical restriction theorems for the Fourier thansform.}
Illiniis Journal of mathematics, Volume 36, Number 1, Spring 1992.

\bibitem{Buldygin}
{\bf Buldygin V.V., Kozachenko Yu.V. }  {\it Metric Characterization of Random
Variables and Random Processes.} 1998, Translations of Mathematics Monograph, AMS, v.188.

\bibitem{Carneiro}
{\bf Emanuel Carneiro, Giuseppe Negro, and Diogo Oliveira E Silva.}
{\it Stability of sharp Fourier restriction to spheres. } \\
arXiv:2108.03412v1  [math.CA]  7 Aug 2021

\bibitem{Ermakov etc. 1986}
{\bf S. V. Ermakov, and E. I. Ostrovsky.} {\it Continuity Conditions, Exponential Estimates, and the Central Limit Theorem for Random Fields.}
 Moscow, VINITY,  1986. (in Russian).

\bibitem{Ferreyra}
{\bf Ferreyra E., Godoy T., Urciuolo M.}  {\it Restriction theorems for the Fourier transform to homogeneous polynomial
surfaces in}  $ \  R^3. \ $  Studia Math. 160, 249 \ - \ 265, 2004.

\bibitem{Fiorenza 1}
{\bf A. Fiorenza.}   {\it Duality and reflexivity in grand Lebesgue spaces. } Collect. Math.
{\bf 51,}  (2000), 131  - 148. \\


\bibitem{Fiorenza 2}
{\bf  4. A. Fiorenza and G.E. Karadzhov.} {\it Grand and small Lebesgue spaces and
their analogs.} Consiglio Nationale Delle Ricerche, Instituto per le Applicazioni
del Calcoto Mauro Picone, Sezione di Napoli, Rapporto tecnico 272/03, (2005).


\bibitem{Fiorenza-Formica-Gogatishvili-DEA2018}
{\bf A.~Fiorenza, M.~R.~Formica} and {\bf A.~Gogatishvili.} {\it On
grand and small Lebesgue and Sobolev spaces and some applications to
PDE's}. \emph{Differ. Equ. Appl.} \textbf{10} (2018), no.~1, 21--46.

\bibitem{fioforgogakoparakoNAtoappear}
{\bf A.~Fiorenza, M. R.~Formica, A.~Gogatishvili, T.~Kopaliani} and
{\bf J.~M. Rakotoson.} {\it Characterization of interpolation
between grand, small or classical Lebesgue spaces}. Preprint
arXiv:1709.05892, Nonlinear Anal., {to appear}.

\bibitem{fioformicarakodie2017}
{\bf A.~Fiorenza, M.~R.~Formica} and {\bf J.~M. Rakotoson.} {\it
Pointwise estimates for {$G\Gamma$}-functions and applications}.
Differential Integral Equations \textbf{30} (2017), no.~11-12,
809--824.

\bibitem{formicagiovamjom2015}
{\bf M.~R. Formica} and {\bf R.~Giova.} {\it Boyd indices in
generalized grand Lebesgue spaces and applications}. Mediterr. J.
Math. \textbf{12} (2015), no.~3, 987--995.

\bibitem{Foschi 1}
{\bf D.Foschi.} {\it Global maximizers for the sphere adjoint Fourier restriction inequality.} J. Funct.
Anal.,  {\bf 268,}  (2015), 690  \ - \  702.

\bibitem{Foschi 2}
{\bf D. Foschi and D. Oliveira e Silva.} {\it Some recent progress in sharp Fourier restriction
theory.}  Preprint, 2017.\\
 arXiv:1701.06895. To appear in Analysis Mathematica.

\bibitem{Frank}
{\bf R. Frank, E. H. Lieb and J. Sabin.}  {\it Maximizers for the Stein - Tomas inequality.} Geom.
Funct. Anal., {\bf 26,}  (2016), no. 4, 1095 m \ - \  1134.

\bibitem{Iwaniec}
{\bf T.Iwaniec and C.Sbordone.} {\it On the integrability of the Jacobian under minimal
hypotheses. } Arch. Rat.Mech. Anal., 119, (1992), 129-143.

\bibitem{Kozachenko 1}
{\bf Kozachenko Yu. V., Ostrovsky E.I. }  (1985). {\it The Banach Spaces of random Variables of subgaussian Type. } Theory of Probab.
and Math. Stat. (in Russian). Kiev, KSU, 32, 43-57. \\

\bibitem{Kozachenko 2}
{\bf 7. Kozachenko Yu.V., Ostrovsky E., Sirota L.}  {\it Relations between exponential tails, moments and
moment generating functions for random variables and vectors.} \\
arXiv:1701.01901v1 [math.FA] 8 Jan 2017 \\

\bibitem{Laura}
{\bf Laura De Carli, Dmitriy Gorbachev, and Sergey Tikhonov.}
{\it Pitt inequalities and restriction theorems for the Fourier transform.}

\bibitem{Ostrovsky 1}
{\bf E.Ostrovsky, L.Sirota, and E.Rogover.} {\it Integral operators in Grand Lebesgue Sraces.}\\
arXiv:0912.2538v1 [math.FA] 13 Dec 2009

\bibitem{Ostrovsky 2}
{\bf E.Ostrovsky, L.Sirota.} {\it Boundedness of operators in bilateral Grand Lebesgue Spaces,} \\
{\it with exact and weakly exact constant calculation.} \\
arXiv:1104.2963v1 [math.FA] 15 Apr 2011

\bibitem{Oliveira 1}
{\bf Diogo Oliveira e Silva.}  {\it A sharp inequality in Fourier restriction theory.}
School of Mathematics, University of Birmingham, Edgbaston, Birmingham B15 2TT, (2014), UK.

\bibitem{Oliveira 2}
{\bf  Diogo Gaspar Teixeira de Oliveira e Silva.}  {\it Sharp Fourier Restriction Theory.}
Mathematisch \ - \ Naturwissenschaftlichen Fakult\"at der Rheinischen Friedrich-Wilhelms-Universit\"at,
Bonn, Mai 2017

\bibitem{Ostrovsky 0}
{\bf Ostrovsky E.I. } (1999). {\it Exponential estimations for Random Fields and its
applications,} (in Russian). Moscow \ - \ Obninsk, OINPE.

\bibitem{Ostrovsky 1}
{\bf Ostrovsky E. and Sirota L. } {\it Sharp moment estimates for polynomial martingales. } \\
arXiv:1410.0739v1 [math.PR] 3 Oct 2014

\bibitem{Ostrovsky 2}
{\bf  Ostrovsky E. and Sirota L. } {\it   Entropy and Grand Lebesgue Spaces approach for the problem  of
Prokhorov - Skorokhod continuity of discontinuous random fields. }\\
arXiv:1512.01909v1 [math.Pr] 7 Dec 2015

\bibitem{Ostrovsky 3}
{\bf  Ostrovsky E. and Sirota L. } {\it  Fundamental function for Grand Lebesgue Spaces.  }
arXiv:1509.03644v1  [math.FA]  11 Sep 2015

\bibitem{Tomas}
{\bf P.A.Tomas.}  {\it A restriction inequality for the Fourier transform. }  Bull. Am. Math. Soc.,
{\bf 81,}  (1975), no. 2, 477 \ - \ 478.

\bibitem{Tao 1}
{\bf T.Tao.}  {\it A sharp bilinear restrictions estimate for paraboloids.} Geom. Funct. Anal.,
{\bf 13,}  (2003), no. 6, 1359 \ - \ 1384.


\bibitem{Tao 2}
{\bf Tao T.}  {\it The Bochner \ - \ Riesz conjecture implies the Restriction conjecture.}  Duke Math. J.,  {\bf 96,}
363 \ - \ 376, 1999.


\bibitem{Tao 3}
{\bf Tao T.}  {\it Some recent progress on the restriction conjecture Fourier Analysis and convexity,}
p.217  \ - \ 243, Appl. Numer. Harmon. Anal., (2004).


\bibitem{Urciuolo}
{\bf Marta Urciuolo.} {\it Restriction of the Fourier transform.} Revista de la Union Matematica Argentina.
Volumen 49, Nrumero 2, 2008, P´aginas 39 \ - \ 44.

\bibitem{Watson}
{\bf G.N.Watson.}  {\it A Treatise on the Theory of Bessel Functions,}  (2nd ed.) Cambridge
University Press, 1966


\end{thebibliography}

\documentstyle[12pt]{article}

  \newcommand{\const}{\rm const}

  \newcommand{\argmax}{\rm  argmax}
  \newcommand{\Rad}{\rm Rad}

\begin{document}

 \begin{center}

 \ {\bf Sharp bounds for  the spherical restriction Fourier transform}\par

\vspace{4mm}

\ {\bf in classical  Lebesgue - Riesz and Grand Lebesgue spaces}\par

\vspace{4mm}

{\bf for  ordinary and radial functions.} \par

\vspace{5mm}

 {\bf M.R.Formica,  E.Ostrovsky and L.Sirota.}

 \end{center}

\vspace{4mm}

 \ Universit\`{a} degli Studi di Napoli Parthenope, via Generale Parisi 13, Palazzo Pacanowsky, 80132,
Napoli, Italy. \\

e-mail: mara.formica@uniparthenope.it \\

\vspace{4mm}

 \ Israel,  Bar - Ilan University, department  of Mathematic and Statistics, 59200, \\

\vspace{4mm}

e-mails: \  eugostrovsky@list.ru \\
sirota3@bezeqint.net \\

\vspace{5mm}

 \begin{center}

 \ {\sc Abstract.}

 \end{center}

\vspace{4mm}

  \ We derive bilateral  estimates for the constants appearing in the Fourier transform restricted  theorems on the Euclidean
  sphere for the ordinary and especially {\it radial} functions  belonging to  the Lebesgue - Riesz spaces as well as belonging
  to the Grand Lebesgue Spaces.  \par
  \ We obtain an exact estimate for the norm  of the restriction Fourier transform operator acting on the radial functions. \par

 \vspace{5mm}

 \ {\it Key words and phrases.} Ordinary and restricted Fourier transform, norm of vector and operator, inner (scalar) product,
  Ordinary and Grand Lebesgue Spaces  (GLS) and norms,  Gaussian function, radii and radial function, Bessel's functions, surface measure,
upper and lower estimates, Tomas - Stein inequality, unitary dilation operator,  extremal function,  examples. \par

\vspace{5mm}

\section{Definitions. Statement of problem. Previous results.}

\vspace{5mm}

\begin{center}

{\sc Fourier transform  and restriction problem .} \\

\end{center}

 \ Define as ordinary the Fourier transform $ \ \hat{f}(k) = \hat{f}[f](k), \ k \in R^d \ $
 for the integrable function on the whole Euclidean space $ \ R^d, \ d = 2,3,\ldots \ $  as follows

\begin{equation} \label{Fourier}
\hat{f}(k) = \hat{f}[f](k) \stackrel{def}{=} \int_{R^d} e^{-i (k,x)} \ f(x) \ dx.
\end{equation}

 \ Henceforth $ \ (k,x) $ denotes as usually the inner (scalar) product  $ \ k x = \sum_{j=1}^d k_j x_j,  \  r = |x|: =
  \sqrt{x,x}, \ s = |k| := \sqrt{k,k}. \ $ \par

\vspace{4mm}

 \ Denote as ordinary by $ \ ||f||_p = ||f||L_p \ $ the classical Lebesgue - Riesz norm for the  measurable function $ \ f: R^d \to R \ $
 or $ \ f: R^d \to C, \ \ C $ is complex plane,  over the whole space

$$
||f||_p = ||f||L_p(R^d) := \left[ \ \int_{R^d} |f(x)|^p \ dx \ \right]^{1/p}.
$$
 \ Denote also the usually {\it surface }  Euclidean measure defined on the measurable subsets of the unit sphere

$$
  S(d-1)  := \{ x, \ x \in R^d, \ |x| = 1 \},
$$
by $ \ \gamma. \ $ For instance,

$$
A(d) := \gamma(S(d-1)) = \frac{ d \ \pi^{d/2}}{\Gamma(1 + d/2)} = \frac{2 \ \pi^{d/2}}{\Gamma(d/2)}
$$
is the area of sphere $ \ S(d-1). \ $ As usually,  $ \ \Gamma(\cdot) \ $ denotes the Euler's Gamma function. \par
 \ The ordinary Lebesgue - Riesz surface norm
 $ \ ||g||L_q(S(d-1)), \ q \ge 1 \ $ for the  measurable function $ \ g: \ S(d-1) \to C \ $ is defined alike one on the whole space

 \begin{equation} \label{surface norm}
 ||g||L_q(S(d-1)) \stackrel{def}{=} \left[ \ \int_{S(d-1)} |g(y)|^q \ \gamma(dy) \ \right]^{1/q}, \ q \ge 1.
 \end{equation}

\vspace{5mm}

 \hspace{2mm} {\bf Definition 1.1. } The inequality of the form

\begin{equation} \label{key ineq}
||\hat{f}||L_q(S(d-1)) \le K_d(p,q) \ ||f||L_p(R^d),
\end{equation}
with finite coefficient $ \ K_d(p,q) \ $ which holds true for arbitrary $ \ f \in L_p(R^d) \ $ and for some non - trivial domain of values $ \ (p,q) \in B, \ $
is named {\it Fourier restriction inequality.}\par
 \ Here the  non - trivial domain  $ \ B = B(d)  \ $ consists on the whole set of the values of the parameters $ \ (p,q), \ p,q \ge 1,   \  $ for which
$ \ K_d(p,q) < \infty. \ $ \par

 \ Another name: Tomas - Stein inequality, see an pioneer work  \cite{Tomas}, 1975; in this article was obtained  in particular the following important necessary
conditions for  this inequality

\begin{equation} \label{necess cond}
1 \le p \le  \frac{2d +2}{d + 3}, \hspace{3mm} q \le \frac{d-1}{d+1} \cdot \frac{p}{p-1}.
\end{equation}

\vspace{4mm}

 \ There are huge works devoted to the calculation, or at last evaluation, of the domain $ \ B, \ $ as well as the  {\it upper} estimation the coefficient $ \ K_d(p,q); \ $
on the other words, the norm of correspondent restriction  Fourier operator on some regular surfaces, in particular on the sphere,  see e.g. in the works
\cite{Bloom}, \cite{Ferreyra}, \cite{Foschi 1} - \cite{Foschi 2}, \cite{Frank}, \cite{Laura}, \cite{Oliveira 1}  - \cite{Oliveira 2}, \cite{Tomas}, \cite{Tao 1} - \cite{Tao 3},
\cite{Urciuolo}  etc.\par

\vspace{4mm}

 \hspace{3mm} {\bf We intent in this short report to offer the simple lower bounds for this coefficients,   calculate its exact value only for the radial functions,
 find in this case the extremal functions. } \par

\vspace{3mm}

  \ {\bf We extent also these results on the Grand Lebesgue Spaces instead ordinary Lebesgue - Riesz ones.} \par

\vspace{4mm}

 \  We can and will to take as the value $ \ K_d(p,q)  \ $ its minimal value:

\begin{equation} \label{def Z}
K_d(p,q) \stackrel{def}{=} \sup_{ f: ||f||_p = 1}  Z_d(||\hat{f}||L_q(S(d-1) ), ||f||_p),
\end{equation}
where

\begin{equation} \label{minimal value}
Z_d(||\hat{f}||L_q(S(d-1) ), ||f||_p) = Z_d(f) \stackrel{def}{=} \ \frac{||\hat{f}||L_q(S(d-1))}{||f||_p}, \ (p,q) \in B,
\end{equation}

\vspace{3mm}

 and separately only for radial functions

\vspace{3mm}

\begin{equation} \label{def Z Ead}
K_d^{Rad}(p,q) \stackrel{def}{=} \sup_{ f: \ f \in \Rad, \ ||f||_p = 1}  Z_d(||\hat{f}||L_q(S(d-1) ), ||f||_p),
\end{equation}

\vspace{4mm}

 \ Introduce also an extremal function,   which is not uniquely determined, e.g. up to change of sign and up to permutation
 of arguments

\begin{equation} \label{extremal fun}
g(x)  = g_d[p,q](x) \stackrel{def}{=} \argmax_{g: \ ||g||_p \le 1}  \ Z_d(||\hat{g}||L_q(S(d-1) ), ||g||_p),
\end{equation}
so that evidently $ \ ||g||_p =  ||g_d[p,q]||_p  = 1 \ $ and

$$
 Z_d(||\hat{g}||L_q(S(d-1) ), ||g||_p) = K_d(p,q), \ (p,q) \in B,
$$
and analogously for the radial functions:

\begin{equation} \label{extremal rad fun}
g^{\Rad}(x)  = g_d^{\Rad}[p,q](x) \stackrel{def}{=} \argmax_{g: \ g \in \Rad, \ ||g||_p \le 1}  \ Z_d(||\hat{g}||L_q(S(d-1) ), ||g||_p),
\end{equation}
so that evidently $ \ ||g^{\Rad}||_p =  ||g_d^{\Rad}[p,q]||_p  = 1 \ $ and of course

$$
 Z_d^{\Rad}(||\hat{g}^{\Rad}||L_q(S(d-1) ), ||g||^{\Rad}_p) = K_d^{\Rad}(p,q), \ (p,q) \in B.
$$

\vspace{5mm}

\begin{center}

 {\sc About radial functions.   }

\end{center}

\vspace{5mm}

 \ Recall that the function $ \ f = f(x), \ x \in R^d \ $ is said to be {\it radial,} or equally spherical symmetry,
 iff it dependent only  on the polar radii (Euclidean norm) of the argument vector  $ \ x: \ r = |x|: \ $

\begin{equation} \label{radial function}
\exists F = F(r), \ f(x) = F(r) = F(|x|), \ x \in R^d; \ r \in (0,\infty).
\end{equation}

 \ Write: $ \  f(\cdot) \in \Rad. \ $ If this function $ \ f(\cdot) \ $ is radial, then   $ \ \hat{f}[f](k) \ $ is one, as well. \par
 \ More detail, introduce the following kernel function

\begin{equation} \label{kernel fun}
V_d(s,r) \stackrel{def}{=} (2 \pi)^{d/2} \ J_{(d-2)/2}( s \ r) \ s^{(2 - d)/2} \ r^{d/2},
\end{equation}
where  $ \ J_l(\cdot) \ $ denotes as usually the Bessel's function  of order $ \ l; \ l \ge 0. \ $\par
 \ Then the function $ \ \hat{f}(k) \ $ has a form $ \ \hat{f}(k) = G(s) = G(|k|), \ $ where

\begin{equation} \label{Fourier transform}
G(s) = \int_0^{\infty} V_d(s, r) \ F(r) \ dr.
\end{equation}

 \vspace{4mm}

 \ In particular, in this radial case

\begin{equation} \label{full int}
\int_{R^d} f(k) \ dk = \int_{R^d} F(|x|) \ dx = G(0+)  = \frac{2}{\Gamma(d/2)} \int_0^{\infty} r^{d-1} \ F(r) \ dr.
\end{equation}

and

\begin{equation} \label{Lp radial norm}
\int_{R^d} |F(|x|)|^p \ dx = \frac{2}{\Gamma(d/2)} \ \int_0^{\infty} r^{d-1} \ |F(r)|^p \ dr.
\end{equation}

\vspace{5mm}

\section{Main result: lower estimate.}

\vspace{5mm}

 \hspace{3mm}  Let us consider the following example (Gaussian radial density function)

\begin{equation} \label{Gaussian density}
 h(x) = h_{\sigma}(x) \stackrel{def}{=} (2 \pi)^{-d/2} \ \sigma^{-d} \ \exp \left\{ \ - 0.5 \ \sigma^{-2} \ ||x||^2  \ \right\}, \ \sigma = \const \in (0,\infty).
\end{equation}
 \ The Fourier transform of this function has a form

$$
\hat{h}(k) = \exp \left\{ \  - 0.5 \ \sigma^2 \ ||k||^2  \ \right\}, \ k \in R^d.
$$

 \ Note that if $ \  k \in  S(d-1),  \ $ then $ \ \hat{h}(k) = \exp \left\{ \ - 0.5 \ \sigma^2 \ \right\}, \ $ so that this function is constant on the
 surface of the unit sphere $ \ S(d-1) \ $ and following

$$
||\hat{h}||_{q, S(d-1)} = \exp(-\sigma^2/2) \ A^{1/q}(d-1), \ q \ge 1.
$$

\vspace{3mm}

 \ Further,

$$
||h||_p = (2 \pi)^{- 0.5 d (1 - 1/p)} \ \sigma^{-d(1 - 1/p)} \ p^{-d/(2p)}, \ p \ge 1.
$$

\ Therefore, for all the positive values $ \ \sigma \ $

$$
K_d(p,q) \ge e^{-\sigma^2/2} \ A^{1/q}(d-1) \ (2 \pi)^{0.5 d (1 - 1/p)} \ p^{d/(2p)} \ \sigma^{d(1 - 1/p)},
$$
and we conclude after maximization over $ \ \sigma \ $ \par

\vspace{4mm}

{\bf Proposition 2.1.}  We deduce that for all the values of parameters for which

$$
d \ge 2, \ p,q \ge 1 \ \Rightarrow \ K_d(p,q) \ge
$$

 \vspace{3mm}

\begin{equation} \label{lower Gaussian}
 A^{1/q}(d-1) \ (2 \pi)^{0.5 d (1 - 1/p)} \ p^{d/(2p)} \ \times [d(1 - 1/p)]^{0.5 d (1 - 1/p)}.
\end{equation}

 \vspace{5mm}

\section{Main result: sharp  radial estimate.}

\vspace{5mm}

 \hspace{3mm}  Note first of all that there ara some reasons to consider the case of radial functions $ \ f = f(|x|). \ $   Indeed,
 denote by $ \  U \ $ an arbitrary unitary linear operator acting from the space $ \ R^d \ $ into oneself.  It follows immediately
 from the expression for this function  (\ref{extremal fun}) after changes of variables that  $ \ g(Ux) = g(x).  \ $  \par
  \ In detail,  introduce the so - called {\it  unitary dilation operator} $ \ T_U[g](x) \ $ for the function $ \ g: R^d \to R \ $  and for
 arbitrary unitary linear operator  $ \ U: R^d \to R^d \ $

 \begin{equation} \label{unit dilat oper}
 T_U[g](x) \stackrel{def}{=} g(Ux), \ x \in R^d.
 \end{equation}
  \ We have

$$
||T_U[g]||_p^p  = \int_{R^d} |g(Ux)|^p \ dx = \int_{R^d} |g(x)|^p \ dx = ||g||_p^p;
$$
 and quite alike

$$
|| \hat{ T_U g  }||L(q, S(d-1)) = || \hat{ g  }||L(q, S(d-1)).
$$

 \ Following, if the function $ \ g(\cdot), \ ||g||_p \le 1 \ $ is extremal and {\it if she was the only one,}  its unitary  delation $ \ T_U [g] \ $
 is also. Therefore, under this hypothesis (which is false!)

 $$
 T_U[g] = \pm g, \ \Leftrightarrow g(\cdot) \in \Rad.
 $$
 \ We can  and will assume without loss of generality that $ \  T_U[g] = g, \ \Rightarrow g(x) = g_o(|x|) \ $  for some (measurable) numerical valued radial
 function $ \ g_o = g_o(r), \ r > 0. \ $    \par

\vspace{3mm}

 \ Thus, the function  $ \ g(x) \ $ is radial, and one can find the extremal in  (\ref{def Z}), (\ref{minimal value}) only among the set of
 all radial functions: $ \ g(x)  = F(|x|) = F(r), \ r > 0.  \ $  \par

 \ Note that here

 \begin{equation} \label{p root spherical}
||F||_p \  = \sqrt[p]{\frac{2}{\Gamma(d/2)} \ \left[ \ \int_0^{\infty} r^{d - 1} \ |F(r)|^p \ dr  \right] }, \  p > 0.
 \end{equation}

 \ Further, it follows from (\ref{Fourier transform})

\begin{equation} \label{Fourier  transform  unit sphere}
G(s) = \int_0^{\infty} V_d(r) \ F(r) \ dr, \ |s| = 1,
\end{equation}
i.e. this function is constant on the surface of unit sphere $ \ S(d-1). \ $ Obviously,

$$
V_d(r) = (2 \pi)^{d/2} \ J_{(d - 2)/2}(r) \ r^{d/2}, \ r > 0.
$$
 \ Therefore

$$
|| G(\cdot) ||_{q, S(d-1)} = A^{1/q}(d-1) \int_0^{\infty} V_d(r) \ F(r) \ dr.
$$

\vspace{5mm}

 \ Thus, the investigated here problem is reduced to the following one: \par

\vspace{4mm}

\begin{equation} \label{def Z0}
W_{d,p,q}[F] \stackrel{def}{=} \left[ \ 2^{-1/p} \ \Gamma^{1/p}(d/2) \ A^{1/q}(d-1) \ \right] \times
\end{equation}

\begin{equation} \label{def ZZ}
\frac{\int_0^{\infty} V_d(r) \ F(r) \ dr }{ \left[ \ \int_0^{\infty} r^{d-1} |F(r)|^p \ dr  \ \right]^{1/p}} \to \sup / \{F \in L(p, R_+, r^{d-1} dr) \}.
\end{equation}

  \ This problem may be transformed as follows. Introduce the following measure $ \ d \nu = \nu_d(dr) := r^{d-1} \ dr;   \  $ then we get to
the following problem

\begin{equation} \label{key problem}
\Phi[F] = \Phi_{d,p,q}[F] \stackrel{def}{=} \frac{\int_0^{\infty} r^{1 - d} \ V_d(r) \ F(r) \ \nu_d(dr)}{||F||_{p, \nu}}
 \to \sup / \{F \in L(p,\nu) \}.
\end{equation}

 \ Denote as usually $ \ p' = p/(p-1), \ p \in (1,\infty); \   1' := +\infty;   \ \infty' := 1.\ $  If follows immediately from  the H\"older's
inequality that

\begin{equation} \label{main result}
Q[d,p,q] \stackrel{def}{=} \sup_{F \in L(p)} \Phi_{d,p,q}[F] = \left\{ \ \int_0^{\infty} r^{p' (1 - d)} \ V_d^{p'}(r) \ \nu_d(dr) \ \right\}^{1/p'}.
\end{equation}

\vspace{4mm}

 \ Let us introduce the following coefficient

\begin{equation} \label{coeff P}
P = P(d; p,q) \stackrel{def}{=} 2^{-1/p} \ \Gamma^{1/p}(d/2) \ A^{1/q}(d-1) \  (2\pi)^{-d/(2 p')}.
\end{equation}

  \ To summarize.\par

  \vspace{4mm}

 \ {\bf Theorem 3.1.}

\begin{equation} \label{rad value}
K^{\Rad}(d; p,q) = P(d;p,q) \cdot \left\{ \ \int_0^{\infty} r^{ (2 + d(p-2))/( 2(p-1) ) } \ J^{p'}_{(d-2)/2}(r) \ dr \  \right\}^{1/p'}.
\end{equation}

\ Herewith the  arbitrary  extremal function $ \ F_o(r) \ $ has a form

\begin{equation} \label{extremal fun}
F_0(r) = C \cdot [V_d(r)]^{1/(p-1)}, \ p > 1, \ C = \const.
\end{equation}

\vspace{4mm}

 \ Notice that the integral in (\ref{rad value}) convergent under our  assumptions iff

\begin{equation} \label{condit converg}
1 < p < \frac{2d}{d + 1}.
\end{equation}

\ As regards for the properties of Bessel's functions see, e.g. \cite{Watson}. Namely, as $ \ z \to 0+ \ $

$$
J_{\alpha}(z) \asymp C_0(\alpha) \ z^{\alpha},
$$
 and correspondingly as $ \ z \to \infty \ $

$$
J_{\alpha}(z) \asymp   C_{\infty}(\alpha) \ z^{-1/2}, \ \alpha = \const \ge 0.
$$

\vspace{5mm}

\section{Generalization on the Grand Lebesgue Spaces.}

\vspace{5mm}

 \hspace{3mm} Let $ \ (a,b) = \const, \ 1 \le a < b \le \infty, \ $ and let $ \ \psi = \psi(p), \ p \in (a,b) \ $  be
 bounded from below: $ \ \inf_{p \in (a,b) } \psi(p) > 0 \ $ measurable function. The set of all such a functions will be
 denoted by  $ \ \Psi(a,b); \ $ put also

$$
\Psi  := \cup_{(a,b): 1 < a  < b < \infty} \Psi(a,b).
$$

\vspace{4mm}

 \ {\bf Definition  4.1.} The Grand Lebesgue Space $ \ G\psi, \ \psi \in \Psi(a,b) \ $  builded over the set $ \ R^d, \ $ or equally over the sphere
 $ \ S(d-1), \ $ consists by definition on all the integrable functions having  a finite norm

\begin{equation} \label{GLS norm Rd}
||f||G\psi(R^d) \stackrel{def}{=} \sup_{p \in (a,b)} \left\{ \ \frac{||f||_p}{\psi(p)}   \ \right\}.
\end{equation}

\vspace{4mm}

 \ These space was investigated in many works, see e.g. \cite{Buldygin}, \cite{Ermakov etc. 1986},
\cite{Fiorenza 1}, \cite{Fiorenza 2}, \cite{Fiorenza-Formica-Gogatishvili-DEA2018},
\cite{fioforgogakoparakoNAtoappear}, \cite{fioformicarakodie2017}, \cite{formicagiovamjom2015}, \cite{Iwaniec},
\cite{Kozachenko 1}, \cite{Kozachenko 2}, \cite{Ostrovsky 0} - \cite{Ostrovsky 3}. In particular, the belonging of
the function to certain Grand Lebesgue Space $ \ G\psi \ $ is closely related with its tail behavior and  is related
with its moment generating function $ \ \int \exp(\lambda f(x)) \ dx. \ $ \par
 \ Suppose now that the function $ \ f = f(x) \ $ belongs to sone GLS $ \ G\psi(a,b), \ \exists \psi \in \Psi: \ $

$$
||f||G\psi \le \psi(p), \ p \in (a,b).
$$
 \ Define the (cut) set

$$
D = \{ \ p, \ \exists q \ \Rightarrow (p,q) \in B \ \},
$$
and assume its non - triviality: $ \ D  \ne \emptyset. \ $ We have

$$
||\hat{f}||_{q,S(d-1)} \le ||f||G\psi \cdot \psi(p) \cdot K(p,q), \ p \in (a,b), \ q \in D.
$$
 \ Therefore

$$
||\hat{f}||_{q,S(d-1)} \le ||f||G\psi \cdot \psi(p) \cdot \zeta(q), \ q \in D,
$$
where

$$
\zeta(q) := \inf_{p \in D} [ \ \psi(p) \cdot K(p,q)  \  ].
$$

 \vspace{3mm}

 \ To summarize: \par

\vspace{4mm}

 \ {\bf Theorem 4.1.}  We deduce  under formulated restrictions

$$
||\hat{f}||_{G\zeta,S(d-1)} \le 1 \times ||f||G\psi,
$$
where the constant $ \ "1" \ $ is the best possible. \par
 \ The non - improvability for this constant is grounded in particular in  \cite{Ostrovsky 1}.

 \vspace{6mm}

\vspace{0.5cm} \emph{Acknowledgement.} {\footnotesize The first
author has been partially supported by the Gruppo Nazionale per
l'Analisi Matematica, la Probabilit\`a e le loro Applicazioni
(GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM) and by
Universit\`a degli Studi di Napoli Parthenope through the project
\lq\lq sostegno alla Ricerca individuale\rq\rq .\par

\end{document}